OpenStudy (anonymous):

What is the pattern here? (1/3) + (1/6) + (1/9) + (1/12) ....

OpenStudy (anonymous):

the denominator goes up by 3 each term

OpenStudy (anonymous):

What would the common difference be then, because I have to put a number

OpenStudy (anonymous):

there is no common different, you are multiplying, not adding to go from one term to the next

OpenStudy (anonymous):

it isnt a difference if it is addition

OpenStudy (anonymous):

Do youmean the solution to the problem?

OpenStudy (anonymous):

in that case you would find the common denominator and solve.

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }+\frac{ 1 }{ 6 }+\frac{ 1 }{ 9 }+\frac{ 1 }{ 12 }\] \[\frac{ 12 }{ 36 }+\frac{ 6 }{ 36 }+\frac{ 4 }{ 36 }+\frac{ 3 }{ 36 }\] \[\frac{ 25 }{ 36 }\]